OpenStudy (anonymous):

In a music competition, a participant has to score a total of at least 40 points in the first four rounds combined to move on to the fifth and final round. Glenn scored 4 points in the first round. He then went on to score additional points in the second, third, and fourth rounds. In each of those rounds, his score was identical. Which inequality best shows the number of points, p, that Glenn scored in each of the second, third, and fourth rounds if he earned a place in the finals? would it be 4+3p less than are =40?

Directrix (directrix):

Are these the options: 4 + 3p ≤ 40 4 + 3p ≥ 40 4p + 3 ≥ 40 4p + 3 ≤ 40

OpenStudy (anonymous):

yes

Directrix (directrix):

Do you have an idea as to which one is the correct option?

OpenStudy (anonymous):

4+3p less than are =40?

Directrix (directrix):

I was thinking 4 + 3p ≥ 40 but I'm not sure it makes sense. Let me think.

OpenStudy (anonymous):

ok

Directrix (directrix):

Glenn earned a place in the finals so he made it to the fifth round. That means he had scored at least 40 points.

Directrix (directrix):

lst round: 4 2nd p 3rd p 4th p -------- 4 + 3p is the number of points he has.

Directrix (directrix):

And, that must be ≥ 40 So, 4 + 3p ≥ 40

Directrix (directrix):

@steffie84 Do you agree?

OpenStudy (anonymous):

am still working on it

Directrix (directrix):

Did you look at my work I posted?

Directrix (directrix):

Glenn has to get 40 or more points before he can move to the 5th level. Do you agree?

OpenStudy (anonymous):

at least 40 points means you either get 40 points, or score more which is greater than or equal to 40 points

Directrix (directrix):

@Nishant_Garg It is how to get the 3p + 4 that I am trying to explain at this point, not the 40 part.

OpenStudy (anonymous):

oh, he scored 4 in the first, p in the 2nd, p in the 3rd, and p in the 4th round total is 4+p+p+p=4+p(1+1+1)=4+p(3)=4+3p

Directrix (directrix):

@steffie84 4 + 3p ≤ 40 is Incorrect.

Directrix (directrix):

@steffie84 I'll leave you with thread hijacker @Nishant_Garg

OpenStudy (anonymous):

ok